# On the (In)Approximability of the Monitoring Edge Geodetic Set Problem

### v1

@@ Abstract
We study the minimum Monitoring Edge Geodetic Set (MEG-Set) problem introduced in [Foucaud et al., CALDAM’23]: given a graph $G$, we say that an edge is monitored by a pair $u,v$ of vertices if all shortest paths between $u$ and $v$ traverse $e$; the goal of the problem consists in finding a subset $M$ of vertices of $G$ such that each edge of $G$ is monitored by at least one pair of vertices in $M$, and $|M|$ is minimized.
In this paper, we prove that all polynomial-time approximation algorithms for the minimum MEG-Set problem must have an approximation ratio of $\Omega(\log n)$, unless $\mathsf{P}$ = $\mathsf{NP}$. To the best of our knowledge, this is the first non-constant inapproximability result known for this problem. We also strengthen the known $\mathsf{NP}$-hardness of the problem on $2$-apex graphs by showing that the same result holds for $1$-apex graphs. This leaves open the problem of determining whether the problem remains $\mathsf{NP}$-hard on planar (i.e., 00-apex) graphs.
On the positive side, we design an algorithm that computes good approximate solutions for hereditary graph classes that admit efficiently computable balanced separators of truly sublinear size. This immediately results in polynomial-time approximation algorithms achieving an approximation ratio of $O(n^{\frac{1}{4}}\sqrt{\log n})$ on planar graphs, graphs with bounded genus, and $k$-apex graphs with $k=O(n^{\frac{1}{4}})$. On graphs with bounded treewidth, we obtain an approximation ratio of $O(\log^{3/2}n)$ for any constant $\varepsilon>0$. This compares favorably with the best-known approximation algorithm for general graphs, which achieves an approximation ratio of $O(\sqrt{n\log n})$ via a simple reduction to the Set Cover problem.
@@ 1 Introduction
Consider a communication network that might suffer link failures. Such a network can be modeled as an undirected and connected graph $G$, where vertices represent hosts and edges correspond to links between hosts. In order to quickly detect failure events, the designer wishes to equip the hosts of the network with additional probes: probes can communicate with one another and they can monitor the distance between the corresponding nodes in the network (e.g., via a traceroute mechanism). Whenever the distance between a pair of vertices increases, this indicates that some communication link is inoperative, and an alarm can be raised. Observe that, in the above scenario, it may be possible for two probes to be connected by two (or more) edge-disjoint shortest paths, in which case they will not be able to detect an edge failure.
@@
This motivates the following problem: given a network $G$, find the smallest possible set $M$ of vertices to equip with probes while ensuring that each edge $e$ of $G$ has at least one pair of vertices $u,v\in M$ for which all shortest paths between $u$ and $v$ in $G$ traverse $e$. The above problem is known as the minimum Monitoring Edge Geodetic Set (MEG-Set) problem and has been introduced in [9], where the authors focus on providing upper and lower bounds on the size of minimum MEG-sets for both general graphs and special graph classes (trees, cycles, unicyclic graphs, complete graphs, grids, hypercubes, …).
@@
Further bounds on the size of MEG-sets have been given for the Cartesian and strong products of two graphs [12], for other graph products including join and direct products [21], as a function of the graph’s girth and chromatic number [6], for ladder, butterfly, circulant and Benes networks, for convex polytopes [20], and for radix triangular mesh networks and Sierpiński graphs [17]. Moreover, the minimum MEG-Set problem was proven to be $\mathsf{NP}$-hard on general graphs [12], $\mathsf{NP}$-hard on $3$-degenerate $2$-apex graphs [8], and $\mathsf{APX}$-hard on $4$-degenerate graphs [7]. If the Exponential Time Hypothesis [13] holds, then the problem cannot be solved in subexponential time on $3$-degenerate graphs [7]. In the same paper, the authors also provide an exact polynomial-time algorithm for interval graphs and two FPT algorithms: one for general graphs parameterized by the sum of cliquewidth and diameter, and the other for chordal graphs parameterized by treewidth. Polynomial-time algorithms for computing optimal MEG-Sets are also known for distance hereditary graphs, $P_{4}$-sparse graphs, bipartite permutation graphs, and strongly chordal graphs [10].
@@
As far as approximation algorithms are concerned, it has been observed [2, 1, 7] that a simple reduction to the Set Cover problem results in a polynomial-time algorithm returning solutions of size $O(k^{2}\log n)$, where $k$ is the size of a MEG-Set of minimum cardinality, thus achieving an approximation ratio of $O(\min\{k\log n,n/k\})=O(\sqrt{n\log n})$. [^1] To the best of our knowledge, this algorithm was first given and analyzed in the Bachelor’s thesis [2] of one of the authors of this manuscript, and the result was subsequently claimed in a short communication based thereon [1]. The same algorithm, and its analysis, also appear in a later unpublished work [7] by a different group of authors.
@@ 1.1 Our results
We prove that the problem is not approximable within a factor of $o(\log(n))$ if $\mathsf{P}\neq\mathsf{NP}$, and that if the stronger assumption $\mathsf{NP}\subseteq\mathsf{DTIME}(n^{O(\log\log n)})$ holds, then the problem admits no $(\alpha\log n)$-approximation algorithm for any constant $\alpha<\frac{1}{2}$. To the best of our knowledge, these are the first non-constant inapproximability results for MEG-Set.
We also extend the aforementioned $O(\sqrt{n\log n})$-approximation algorithm for MEG-Set to a generalized version of the problem in which edges have non-negative lengths, vertices have binary costs, and a minimum-cost subset of vertices that monitors a given subset of edges needs to be selected. We use this result as a building block to design a more involved approximation algorithm for MEG-Set that relies on recursively computing vertex-separators and provides good approximations for graph families that admit sparse balanced vertex-separators. More precisely, we achieve an approximation ratio of $O(n^{\frac{1}{4}}\sqrt{\log n})$ on planar graphs, on graphs with bounded genus, and on $k$-apex graphs when $k=O(n^{\frac{1}{4}})$. We also obtain an approximation ratio of $O(\log^{3/2}n)$ on graphs with bounded treewidth.
@@ 1.2 Organization of the paper
We start by giving some preliminary definitions and notation in Section 2. Our inapproximability result for general graphs is presented in Section 3, while the $\mathsf{NP}$-Hardness result for $1$-apex graphs is discussed in Section 4. Finally, our approximation algorithms are given in Section 5.
@@ 3 Inapproximability of MEG-Set
We reduce from the Set Cover problem. A Set Cover instance $\mathcal{I}=\langle X,\mathcal{S}\rangle$ is described as a set of $\eta$ items $X=\{x_{1},\dots,x_{\eta}\}$ and a collection $\mathcal{S}=\{S_{1},\dots,S_{h}\}$ of $h\geq 2$ distinct subsets of $X$, such that each subset contains at least two items and each item appears in at least two subsets. [^2] This can be guaranteed w.l.o.g. by repeatedly reducing the instance by applying the first applicable of the following two reduction rules. Rule 1: if there exists an item $x_{i}$ that is contained in a single subset $S_{j}$, then $S_{j}$ belongs to all feasible solutions, and we reduce to the instance in which both $S_{j}$ and $x_{i}$ have been removed. Rule 2: if there exists a subset $S_{j}$ that contains a single element, then (due to Rule 1) there is an optimal solution that does not contain $S_{j}$, and we reduce to the instance in which $S_{j}$ has been removed. Notice that this process can only decrease the values of $\eta$ and $h$. The goal is that of computing a collection $\mathcal{S}^{*}\subseteq\mathcal{S}$ of minimum size such that $\cup_{S_{i}\in\mathcal{S}^{*}}S_{i}=X$ [^3] We assume w.l.o.g. that $\cup_{i=1}^{h}S_{i}=X$, i.e., that a solution exists. It is known that, unless $\mathsf{P}=\mathsf{NP}$, all polynomial-time approximation algorithms for the Set Cover must have an approximation ratio of $(1-o(1))\ln\eta$ when $h=O(\text{poly}(\eta))$ [18, 4]. Moreover, unless $\mathsf{NP}\subseteq\mathsf{DTIME}(n^{O(\log\log n)})$, the same inapproximability result holds even for $h\leq\eta$ [5].
@@
Let $k$ be an integer parameter, whose exact value will be chosen later, that satisfies $2\leq k=O(\text{poly}(N))$. We construct a graph $G$ that contains $k$ copies $H_{1},\ldots,H_{k}$ of $H$ as induced subgraphs. In the following, for any $\ell=1,\ldots,k$, we denote by $x_{i,\ell}$ and $S_{j,\ell}$ the vertices of $H_{\ell}$ corresponding to the vertices $x_{i}$ and $S_{j}$ of $H$, respectively. More precisely, we build $G$ by starting with a graph that contains exactly the $k$ copies $H_{1},\ldots,H_{k}$ of $H$ and augmenting it as follows (see Figure 1):
@@
Let $Y=\{y_{i}\mid i=1,\dots,\eta\}$, $Y^{\prime}=\{y^{\prime}_{i}\mid i=1,\dots,\eta\}$, $Z=\{z_{j}\mid j=1,\dots,h\}$, and $Z^{\prime}=\{z_{j}\mid i=1,\dots,h\}$. Moreover, define $L$ as the set of all vertices of degree $1$ in $G$, i.e., $L=Y^{\prime}\cup Z^{\prime}$. By Lemma 2.1, the vertices in $L$ belong to all MEG-sets of $G$.
@@
Proof 3.4.
Since $L\subseteq M$, by Lemma 3.1, we only need to argue about edges with at least one endvertex in some $H_{\ell}$, with $1\leq\ell\leq k$. Let $S_{j}\in\mathcal{S}_{\ell}$, and consider any $x_{i}\in S_{j}$. Edge $(S_{j,\ell},z_{j})$ is monitored by $\{z^{\prime}_{j},S_{j,\ell}\}$. Edges $(S_{j,\ell},x_{i,\ell})$ and $(x_{i,\ell},y_{i})$ are monitored by $\{S_{j,\ell},y^{\prime}_{i}\}$.
@@
We say that a MEG-set $M$ is minimal if, for every $v\in M$, $M\setminus\{v\}$ is not a MEG-set. Lemma 2.2 ensures that any minimal MEG-set $M$ does not contain any of the vertices $y_{i}$, for $i=1,\dots,\eta$, or $z_{j}$ for $j=1,\dots,h$. Hence, $M\setminus L$ contains only vertices in $\bigcup_{\ell=1}^{k}V(H_{\ell})$.
@@
Proof 3.8.
Let $M^{\prime\prime}$ be a minimal MEG-set of $G$ that is obtained from $M^{\prime}$ by possibly discarding some of the vertices. Clearly, $|M^{\prime\prime}|\leq|M^{\prime}|$ and $M^{\prime\prime}$ can be computed in polynomial time. Moreover, by Lemma 3.5, for every $i=1,\ldots,\eta$ and every $\ell=1,\ldots,k$, $M^{\prime\prime}$ contains $x_{i,\ell}$ or some $S_{j,\ell}$ such that $S_{j}$ covers $x_{i}$. We compute $M$ from $M^{\prime\prime}$ by replacing each $x_{i,\ell}\in M^{\prime\prime}$ with $S_{j,\ell}$, where $S_{j}\in\mathcal{S}$ is any set that covers $x_{i}$. As a consequence, for every $\ell=1,\ldots,k$, the set $\mathcal{S}_{\ell}=\{S_{j}\in\mathcal{S}\mid S_{j,\ell}\in M\}$ is a set cover of $\mathcal{I}$. Moreover, since $M^{\prime\prime}$ contains all vertices in $L$ by Lemma 2.1, so does $M$. Then, Lemma 3.3 implies that $M$ is a MEG-set of $G$.
@@
We now construct the graph $G$ with $n=(k+2)N\leq N^{2}$ vertices by making $k=N-2$ copies of $H$. Next, we run the $(\alpha\ln n)$-approximation algorithm to compute a MEG-set $M^{\prime}$ of $G$, and we use Lemma 3.7 to find a MEG-set $M$ with $|M|\leq|M^{\prime}|$ that contains $k$ set covers $\mathcal{S}_{1},\ldots,\mathcal{S}_{k}$ in polynomial time. Among these $k$ set covers, we output one $\mathcal{S}^{\prime}$ of minimum size.
To analyze the approximation ratio of the above algorithm, let $M^{*}$ be an optimal MEG-set of $G$. Lemma 3.3 ensures that $|M^{*}|\leq|L|+kh^{*}=N+kh^{*}$, and hence:
@@
It is known that, unless $\mathsf{P}=\mathsf{NP}$, there exists a constant $c\geq 1$ such that all polynomial-time approximation algorithms for instances of Set Cover having $\eta$ items and $h\leq\eta^{c}$ sets must have an approximation factor of $(1-o(1))\ln\eta$ [18, 4]. Then, Lemma 3.9 with any constant $\alpha<\frac{1}{2c}$ and $\varepsilon=(\frac{1}{2}-\alpha c)^{2}$ implies that no polynomial-time $(\alpha\ln n)$-approximation algorithm for MEG-Set can exist, unless $\mathsf{P}=\mathsf{NP}$.
Similarly, assuming $\mathsf{NP}\subseteq\mathsf{DTIME}(n^{O(\log\log n)})$, we can choose $c=1$ [5] to rule out any polynomial-time algorithm with an approximation of $\alpha\ln n$ for MEG-Set, where $\alpha$ is a constant strictly smaller than $\frac{1}{2}$.

### v3

@@ Abstract
We study the minimum Monitoring Edge Geodetic Set (MEG-Set) problem introduced in [Foucaud et al., CALDAM’23]: given a graph $G$, we say that an edge is monitored by a pair $u,v$ of vertices if all shortest paths between $u$ and $v$ traverse $e$; the goal is to find a subset $M$ of vertices of $G$ such that each edge of $G$ is monitored by at least one pair of vertices in $M$, and $|M|$ is minimized.
In this paper, we prove that all polynomial-time approximation algorithms for the minimum MEG-Set problem must have an approximation ratio of $\Omega(\log n)$, unless $\mathsf{P}$ = $\mathsf{NP}$. To the best of our knowledge, this is the first non-constant inapproximability result known for this problem. We also strengthen the known $\mathsf{NP}$-hardness of the problem on $2$-apex graphs by showing that the same result holds for $1$-apex graphs. This leaves open the question of determining whether the problem remains $\mathsf{NP}$-hard on planar (i.e., 0-apex) graphs.
On the positive side, we design an algorithm that computes good approximate solutions for hereditary graph classes that admit efficiently computable balanced separators of truly sublinear size. This immediately yields polynomial-time approximation algorithms achieving an approximation ratio of $O(n^{\frac{1}{4}}\sqrt{\log n})$ on planar graphs, graphs with bounded genus, and $k$-apex graphs with $k=O(n^{\frac{1}{4}})$. On graphs with bounded treewidth, we obtain an approximation ratio of $O(\log^{3/2}n)$. This compares favorably with the best-known approximation algorithm for general graphs, which achieves an approximation ratio of $O(\sqrt{n\log n})$ via a simple reduction to the Set Cover problem.
@@ 1 Introduction
Consider a communication network that might suffer link failures. Such a network can be modeled as an undirected and connected graph $G$, where vertices represent hosts and edges correspond to links between hosts. In order to quickly detect failure events, the designer wishes to equip the hosts of the network with additional probes: probes can communicate with one another and they can monitor the distance between their corresponding nodes in the network (e.g., via a traceroute mechanism). Whenever the distance between a pair of vertices increases, this indicates that some communication link is inoperative, and an alarm can be raised. Observe that, in the above scenario, it may be possible for two probes to be connected by two (or more) edge-disjoint shortest paths, in which case they will not be able to detect an edge failure.
@@
This motivates the following problem: given a network $G$, find the smallest possible set $M$ of vertices to equip with probes while ensuring that each edge $e$ of $G$ has at least one pair of vertices $u,v\in M$ for which all shortest paths between $u$ and $v$ in $G$ traverse $e$. The above problem is known as the minimum Monitoring Edge Geodetic Set (MEG-Set) problem and was introduced in [9], where the authors focus on providing upper and lower bounds on the size of minimum MEG-Sets for both general graphs and special graph classes (trees, cycles, unicyclic graphs, complete graphs, grids, hypercubes, …).
@@
Further bounds on the size of MEG-Sets have been given for the Cartesian and strong products of two graphs [12], for other graph products including join and direct products [21], as a function of the graph’s girth and chromatic number [6], for ladder, butterfly, circulant and Benes networks, for convex polytopes [20], and for radix triangular mesh networks and Sierpiński graphs [17]. Moreover, the minimum MEG-Set problem was proven to be $\mathsf{NP}$-hard on general graphs [12], $\mathsf{NP}$-hard on $3$-degenerate 22-apex graphs [8], and $\mathsf{APX}$-hard on $4$-degenerate graphs [7]. If the Exponential Time Hypothesis [13] holds, then the problem cannot be solved in subexponential time on $3$-degenerate graphs [7]. In the same paper, the authors also provide an exact polynomial-time algorithm for interval graphs and two FPT algorithms: one for general graphs parameterized by the sum of cliquewidth and diameter, and the other for chordal graphs parameterized by treewidth. Polynomial-time algorithms for computing optimal MEG-Sets are also known for distance-hereditary graphs, $P_{4}$-sparse graphs, bipartite permutation graphs, and strongly chordal graphs [10].
@@
As far as approximation algorithms are concerned, it has been observed [2, 1, 7] that a simple reduction to the Set Cover problem yields a polynomial-time algorithm returning solutions of size $O(k^{2}\log n)$, where $k$ is the size of a MEG-Set of minimum cardinality, thus achieving an approximation ratio of $O(\min\{k\log n,n/k\})=O(\sqrt{n\log n})$. [^1] To the best of our knowledge, this algorithm was first given and analyzed in a Bachelor’s thesis [2], and the result was subsequently claimed in a short communication based thereon [1]. The same algorithm, and its analysis, also appear in a later unpublished work [7] by a different group of authors.
@@ 1.1 Our results
We prove that, if $\mathsf{P}\neq\mathsf{NP}$, the problem admits no polynomial-time approximation algorithm with an approximation factor of $o(\log n)$. Moreover, the problem admits no $(\alpha\log n)$-approximation algorithm for any constant $\alpha<\frac{1}{2}$ unless $\mathsf{NP}\subseteq\mathsf{DTIME}(n^{O(\log\log n)})$. To the best of our knowledge, these are the first non-constant inapproximability results for the minimum MEG-Set problem.
We also extend the aforementioned $O(\sqrt{n\log n})$-approximation algorithm for minimum MEG-Set to a generalized version of the problem in which edges have non-negative lengths, vertices have binary costs, and a minimum-cost subset of vertices that monitors a given subset of edges needs to be selected. We use this result as a building block to design a more involved approximation algorithm for minimum MEG-Set that relies on recursively computing vertex-separators and provides good approximations for graph families that admit sparse balanced vertex-separators. More precisely, we achieve an approximation ratio of $O(n^{\frac{1}{4}}\sqrt{\log n})$ on planar graphs, on graphs with bounded genus, and on $k$-apex graphs when $k=O(n^{\frac{1}{4}})$. We also obtain an approximation ratio of $O(\log^{3/2}n)$ on graphs with bounded treewidth.
@@ 1.2 Organization of the paper
We start by giving some preliminary definitions and notation in Section˜2. Our inapproximability result for general graphs is presented in Section˜4, while the $\mathsf{NP}$-Hardness result for $1$-apex graphs is given in Section˜3. Finally, Section˜5 is devoted to our approximation algorithms.
@@ 2 Preliminaries and notation
3 $\mathsf{NP}$-Hardness of minimum MEG-Set on graphs with apex 11
In this section, we show a reduction from the minimum Planar Dominating Set problem on graphs with girth at least seven (which is known to be $\mathsf{NP}$-Hard [22]) to the minimum MEG-Set problem on graphs with apex $1$.
Let $G=(V,E)$ be a planar input graph with $|V(G)|\geq 2$ and with girth $g(G)\geq 7$. We construct an instance of MEG-Set, namely $H$, starting from $G$ and augmenting it as follows: for each vertex $v\in V(G)$, we add 2 vertices $v^{\prime}$ and $v^{\prime\prime}$ to $G$, then connect $v$ to $v^{\prime}$ and $v^{\prime}$ to $v^{\prime\prime}$. We denote the set containing all the vertices $v^{\prime}$ as $L^{\prime}$ and the set containing all the vertices $v^{\prime\prime}$ as $L^{\prime\prime}$. Notice that:
  1. i)
  2. ii)
The basic idea is to monitor each edge that is incident to a vertex $v\in V(G)$ that belongs to a selected MEG-Set in $H$. We denote the set of edges $(v,v^{\prime})$ as $E_{L^{\prime}}$ and the set of edges $(v^{\prime},v^{\prime\prime})$ as $E_{L^{\prime\prime}}$. Note that at this point of the construction, all edges are monitored by at least one pair of leaves in $L^{\prime\prime}$. To avoid this, we shortcut all the shortest paths from one vertex in $L^{\prime\prime}$ to another by inserting two vertices $v_{*}$ and $v_{*}^{\prime}$ and connecting them together. Then, we connect $v_{*}$ to all vertices $v^{\prime}\in L^{\prime}$. We denote the set containing edges $(v_{*},v^{\prime})$ with $v^{\prime}\in L^{\prime}$ as $E_{*}$. The vertex $v_{*}$ is used to create shortcuts between each shortest path from a vertex in $L^{\prime}\cup L^{\prime\prime}$ to a vertex in $L^{\prime}\cup L^{\prime\prime}$. Notice that, since $v_{*}^{\prime}$ is a leaf, it belongs to every MEG-Set of $G$ due to Lemma˜2.1. The vertex $v_{*}^{\prime}$ is used to monitor the edges in $E_{*}$. Formally, let $H=(V(G)\cup L^{\prime}\cup L^{\prime\prime}\cup\{v_{*},v_{*}^{\prime}\},E(G)\cup E_{L^{\prime}}\cup E_{L^{\prime\prime}}\cup E_{*}\cup\{\;\{v_{*},v_{*}^{\prime}\}\;\})$, where:
  • $L^{\prime}=\{v^{\prime}:v\in V(G)\}$;
  • $L^{\prime\prime}=\{v^{\prime\prime}:v\in V(G)\}$;
  • $E_{L^{\prime}}=\{\{v,v^{\prime}\}:v\in V(G)\}$;
  • $E_{L^{\prime\prime}}=\{\{v^{\prime\prime},v^{\prime}\}:v\in V(G)\}$;
  • $E_{*}=\{\{v_{*},v^{\prime}\}:v\in V(G)\}$.
$a^{\prime}bb^{\prime}cc^{\prime}dd^{\prime}a′′a^{\prime\prime}$$b^{\prime\prime}$$c^{\prime\prime}$$d^{\prime\prime}$$_{*}vv_{*}^{\prime}$
Figure 1: An example of the construction of the graph $H$ from the input graph $G=(\{a,b,c,d\},\{\;\{a,b\},\{b,c\},\{c,d\},\{d,a\}\;\})$, which is shown in black.
We prove the $\mathsf{NP}$-Completeness of minimum MEG-Set on graphs with apex $1$ via two lemmas. The first lemma shows that if there is a MEG-Set of the graph $H$ having size at most $k+|V(G)|+1$, there exists a Dominating Set of $G$ having size at most $k$. Conversely, the second lemma shows that if there is a Dominating Set of the graph $G$ having size at most $k$, there exists a MEG-Set of $H$ having size $k+|V(G)|+1$.
Lemma 3.1.
Proof 3.2.
  • Case 3: $e\in E_{L^{\prime}}$. Let $e=(u,u^{\prime})$ where $u$ is the unique endpoint of $e$ in $V(G)$. Since $D$ dominates all the vertices in $V(G)$, either $u\in D$, and we let $x=u$, or $u\not\in D$ and there exists some vertex $x\in D\cap N(u)$. There is a single shortest path from $u^{\prime\prime}$ to $x$, and it traverses the edge $(u,u^{\prime})$.
Lemma 3.3.
Proof 3.4.
Theorem 3.5.
Proof 3.6.
For each planar graph $G$ with girth at least seven, it is possible to construct in polynomial time a graph $H$ with apex $1$, as discussed in this section. From Lemmas 3.1 and 3.3, there exists a Dominating Set $D$ of $G$ such that $|D|\leq k$ if and only if there exists a MEG-Set $M$ of $H$ such that $|M|\leq k+|V(G)|+1$. Since the Dominating Set problem is $\mathsf{NP}$-Hard even for planar graphs with girth at least seven, this implies that the MEG-Set decision problem is $\mathsf{NP}$-Hard on graphs with apex $1$.
@@ 4 Inapproximability of MEG-Set
We reduce from the Set Cover problem. A Set Cover instance $\mathcal{I}=\langle X,\mathcal{S}\rangle$ is described as a set of $\eta$ items $X=\{x_{1},\dots,x_{\eta}\}$ and a collection $\mathcal{S}=\{S_{1},\dots,S_{h}\}$ of $h\geq 2$ distinct subsets of $X$, such that each subset contains at least two items and each item appears in at least two subsets. [^2] This can be guaranteed w.l.o.g. by repeatedly reducing the instance by applying the first applicable of the following two reduction rules. Rule 1: if there exists an item $x_{i}$ that is contained in a single subset $S_{j}$, then $S_{j}$ belongs to all feasible solutions, and we reduce to the instance in which both $S_{j}$ and $x_{i}$ have been removed. Rule 2: if there exists a subset $S_{j}$ that contains a single element, then (due to Rule 1) there is an optimal solution that does not contain $S_{j}$, and we reduce to the instance in which $S_{j}$ has been removed. Notice that this process can only decrease the values of $\eta$ and $h$. The goal is that of computing a collection $\mathcal{S}^{*}\subseteq\mathcal{S}$ of minimum size such that $\cup_{S_{i}\in\mathcal{S}^{*}}S_{i}=X$. [^3] We assume w.l.o.g. that $\cup_{i=1}^{h}S_{i}=X$, i.e., that a solution exists. It is known that, unless $\mathsf{P}=\mathsf{NP}$, all polynomial-time approximation algorithms for the Set Cover must have an approximation ratio of $(1-o(1))\ln\eta$ when $h=O(\text{poly}(\eta))$ [18, 4]. Moreover, unless $\mathsf{NP}\subseteq\mathsf{DTIME}(n^{O(\log\log n)})$, the same inapproximability result holds even for $h\leq\eta$ [5].
@@
Let $k$ be an integer parameter, whose exact value will be chosen later, that satisfies $2\leq k=O(\text{poly}(N))$. We construct a graph $G$ that contains $k$ copies $H_{1},\ldots,H_{k}$ of $H$ as induced subgraphs. In the following, for any $\ell=1,\ldots,k$, we denote by $x_{i,\ell}$ and $S_{j,\ell}$ the vertices of $H_{\ell}$ corresponding to the vertices $x_{i}$ and $S_{j}$ of $H$, respectively. More precisely, we build $G$ by starting with a graph that contains exactly the $k$ copies $H_{1},\ldots,H_{k}$ of $H$ and augmenting it as follows (see Figure˜2):
@@
Let $Y=\{y_{i}\mid i=1,\dots,\eta\}$, $Y^{\prime}=\{y^{\prime}_{i}\mid i=1,\dots,\eta\}$, $Z=\{z_{j}\mid j=1,\dots,h\}$, and $Z^{\prime}=\{z_{j}\mid i=1,\dots,h\}$. Moreover, define $L$ as the set of all vertices of degree $1$ in $G$, i.e., $L=Y^{\prime}\cup Z^{\prime}$. By Lemma˜2.1, the vertices in $L$ belong to all MEG-Sets of $G$.
@@
Proof 4.4.
Since $L\subseteq M$, by Lemma˜4.1, we only need to argue about edges with at least one endvertex in some $H_{\ell}$, with $1\leq\ell\leq k$. Let $S_{j}\in\mathcal{S}_{\ell}$, and consider any $x_{i}\in S_{j}$. Edge $(S_{j,\ell},z_{j})$ is monitored by $\{z^{\prime}_{j},S_{j,\ell}\}$. Edges $(S_{j,\ell},x_{i,\ell})$ and $(x_{i,\ell},y_{i})$ are monitored by $\{S_{j,\ell},y^{\prime}_{i}\}$.
@@
We say that a MEG-Set $M$ is minimal if, for every $v\in M$, $M\setminus\{v\}$ is not a MEG-Set. Lemma˜2.2 ensures that any minimal MEG-Set $M$ does not contain any of the vertices $y_{i}$, for $i=1,\dots,\eta$, or $z_{j}$ for $j=1,\dots,h$. Hence, $M\setminus L$ contains only vertices in $\bigcup_{\ell=1}^{k}V(H_{\ell})$. The following lemma characterizes the structure of minimal MEG-Sets of $G$.
@@
Proof 4.8.
Let $M^{\prime\prime}$ be a minimal MEG-Set of $G$ that is obtained from $M^{\prime}$ by possibly discarding some of the vertices. Clearly, $|M^{\prime\prime}|\leq|M^{\prime}|$ and $M^{\prime\prime}$ can be computed in polynomial time. Moreover, by Lemma˜4.5, for every $i=1,\ldots,\eta$ and every $\ell=1,\ldots,k$, $M^{\prime\prime}$ contains $x_{i,\ell}$ or some $S_{j,\ell}$ such that $S_{j}$ covers $x_{i}$. We compute $M$ from $M^{\prime\prime}$ by replacing each $x_{i,\ell}\in M^{\prime\prime}$ with $S_{j,\ell}$, where $S_{j}\in\mathcal{S}$ is any set that covers $x_{i}$. As a consequence, for every $\ell=1,\ldots,k$, the set $\mathcal{S}_{\ell}=\{S_{j}\in\mathcal{S}\mid S_{j,\ell}\in M\}$ is a set cover of $\mathcal{I}$. Moreover, since $M^{\prime\prime}$ contains all vertices in $L$ by Lemma˜2.1, so does $M$. Then, Lemma˜4.3 implies that $M$ is a MEG-Set of $G$.
@@
We now construct the graph $G$ with $n=(k+2)N\leq N^{2}$ vertices by making k=N2k=N-2 copies of $H$. Next, we run the $(\alpha\ln n)$-approximation algorithm to compute a MEG-Set $M^{\prime}$ of $G$, and we use Lemma˜4.7 to find a MEG-Set $M$ with $|M|\leq|M^{\prime}|$ that contains $k$ set covers $\mathcal{S}_{1},\ldots,\mathcal{S}_{k}$ in polynomial time. Among these $k$ set covers, we output one $\mathcal{S}^{\prime}$ of minimum size.
To analyze the approximation ratio of the above algorithm, let $M^{*}$ be an optimal MEG-Set of $G$. Lemma˜4.3 ensures that $|M^{*}|\leq|L|+kh^{*}=N+kh^{*}$, and hence:
@@
It is known that, unless $\mathsf{P}=\mathsf{NP}$, there exists a constant $c\geq 1$ such that all polynomial-time approximation algorithms for instances of Set Cover having $\eta$ items and $h\leq\eta^{c}$ sets must have an approximation factor of $(1-o(1))\ln\eta$ [18, 4]. Then, Lemma˜4.9 with any constant $\alpha<\frac{1}{2c}$ and $\varepsilon=(\frac{1}{2}-\alpha c)^{2}$ implies that no polynomial-time $(\alpha\ln n)$-approximation algorithm for MEG-Set can exist, unless $\mathsf{P}=\mathsf{NP}$.
Similarly, we can choose $c=1$ [5] to rule out any polynomial-time algorithm with an approximation of $\alpha\ln n$ for MEG-Set, where $\alpha$ is a constant strictly smaller than $\frac{1}{2}$, unless $\mathsf{NP}\subseteq\mathsf{DTIME}(n^{O(\log\log n)})$.